OpenStudy (anonymous):

What is the area of a circle with a radius of 3 meters? Use twenty-two sevenths for pi.

OpenStudy (anonymous):

Area = 22/7*32 = 28.'285714' square meters and the decimals repeated

OpenStudy (calculusxy):

Area of circle=r^2*pi